OpenStudy (nuccioreggie):

An equation is shown below: 6(2x – 11) + 15 = 3x + 12 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points)

OpenStudy (whovianchick):

Part A: 6(2x-11) + 15 = 3x +12 12x-66 + 15 = 3x+12 (distribute 6 times 2x and -11) 12x -51 = 3x+12 (combine -66 and 15) 9x-51 = 12 (subtract 3x from both sides) 9x = 63 (add 51 to both sides) x= 7 (divide both sides by 9) Part B: x=7